OpenStudy (anonymous):

Please Help: lim as x approaches 0 ((sinx)(1-cosx))/(-3x^6)

OpenStudy (anonymous):

=(sin(x)-sin(x)cos(x))/(-3x^6)) =(sin(x)-sin(2x)/2)/(-3x^6). takes the series expansion of sin(x). =((x-x^3/6+O(4)-(x-2x^3/3+O(4)))/(-3x^6) =(x^3/2+O(4))/(-3x^6) =-1/(6x^3)-O(4)/(6x^6) =-infinity.

OpenStudy (anonymous):

what is the series expansion of sinx?

OpenStudy (anonymous):

the series expansion about 0 is x-x^3/6+x^5/120-x^7/5040+...

OpenStudy (anonymous):

is there a way to do it without series expansion? I know we learned the rule of sinx/x = 1 but when I apply that the other section is still 0 over 0

OpenStudy (anonymous):

you can do it by taking two derivatives instead of one

OpenStudy (anonymous):

well I would do that but we have to show all work and for limits at zero we don't use derivatives

OpenStudy (tkhunny):

\(\dfrac{\sin(x)\cdot(1-\cos(x))}{-3x^{6}} = \dfrac{\sin(x)}{x}\cdot\dfrac{1-\cos(x)}{-3x^{5}} = \dfrac{\sin(x)}{x}\cdot\dfrac{1-\cos^{2}(x)}{-3x^{5}\cdot(1+cos(x))}\) It is enough.

OpenStudy (anonymous):

or would I be able to break it up into lim (sinx)/(x^6) + lim (1-cosx)/-3 and get infinity +0

OpenStudy (anonymous):

that works

OpenStudy (anonymous):

really? so that is a legitimate rule?

OpenStudy (anonymous):

actually, the limits would have to be multiplied. nvm.

OpenStudy (anonymous):

what tk said works though

OpenStudy (anonymous):

im following tks but I cant see the whole thing

OpenStudy (tkhunny):

Well, forget that last part. It doesn't get us anywhere. Splitting off the sine simplifies the issue, but doesn't settle the matter.

OpenStudy (tkhunny):

On Third thought \(\dfrac{\sin(x)}{x}\cdot\dfrac{1-\cos^{2}(x)}{-3x^{5}(1+\cos(x))} = \dfrac{\sin(x)}{x}\cdot\dfrac{\sin^{2}(x)}{-3x^{5}(1+\cos(x))}\) \(= \dfrac{\sin(x)}{x}\cdot\dfrac{\sin^{2}(x)}{x^{2}}\cdot\dfrac{1}{-3x^{5}(1+\cos(x))}\) NOW! We have settled the matter. The two sine pieces are one (1). The last piece BLOWS UP!

OpenStudy (tkhunny):

Whoops! That last exponent on x in the denominator should be 3. The result is unaltered.

OpenStudy (anonymous):

so the answer is 1?

OpenStudy (tkhunny):

?? How did you conclude that? "BLOWS UP" is usually not good. There is no limit as x approaches zero (0). If it were either or both of the sine pieces, yes, it would be just one (1).

OpenStudy (anonymous):

so that third piece doesn't exsist causing the rest of the limit to not exsist?

OpenStudy (anonymous):

it exists; it is equal to negative infinity.

OpenStudy (tkhunny):

The other two limits are FINITE. They are easily overcome by one piece that is not finite.

OpenStudy (anonymous):

okay i get it now! thanks guys so much!

OpenStudy (tkhunny):

Infinity is not a place. The limit CANNOT be "infinity" in the Real Numbers. "Infinity" is NOT a Real Number. Besides, in order for a limit to exist, both it's left- and right-hand limits must agree. Not the case on this one. Approach from the positive side and it wanders off without bound in the negative direction. Approach from the negative side and it wanders off without bound in the positive direction. The limit does not exist.